40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7337     Entry time: Tue Sep 4 13:50:26 2012
Author: ericq 
Type: Update 
Category: PSL 
Subject: PMC Realigned, power adjusted 

 I adjusted the PMC alignment this morning, brought the transmission up to 0.83V.

After the lunch meeting, we found the the MC transmission was higher than recently seen. Turned out the HWP had drifted, causing 30mW to be input to the MC. I adjusted it back down to 20mW. 

ELOG V3.1.3-